OpenStudy (anonymous):

if each interior angle of a regular polygon measures 144 degrees, how many sides does the polygon have?

OpenStudy (jdoe0001):



OpenStudy (jdoe0001):



OpenStudy (jdoe0001):

so what do you think angle "x"?

OpenStudy (wolf1728):

Polygon angles = (number of sides -2) * 180 So triangle has 180° total square = 360° Now the measure of EACH angle is given by "Average" angle = [(number of sides -2) * 180] / number of sides so for 144° we have 144° = [(number of sides -2) * 180] / number of sides number of sides = # sides-2 * 180 / 144 (number of sides / number of sides-2) = 180 / 144 I managed to work up this formula but haven't been able to get beyond the "(number of sides / number of sides-2)" step

OpenStudy (wolf1728):

(number of sides / number of sides-2) = 180 / 144 Dividing top and bottom of (180/44) by 36 (number of sides / number of sides-2) = 5 / 4 (n / n-2) = 5 / 4 n = (5n -10) / 4 (5n -10) / n = 4 multiplying both sides by n 5n - 10 = 4 * n 5n -4n = 10 n = 10